Solve the following equations and check your results. 4z + 3 = 6 + 2z 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

4z+3=6+2z
⇒4z-2z=6-3
⇒2z=3
⇒z=3/2
Putting the value of z in RHS and LHS we get,
(4×3/2)+3=6+(2×3/2)
⇒6+3=6+3
⇒9=9
⇒LHS=RHS

Thanks